A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is
20%
salt and Solution B is
45%
salt. She wants to obtain
110
ounces of a mixture that is
25%
salt. How many ounces of each solution should she use?

Answer by stanbon(75887)

She knows that Solution A is
20%
salt and Solution B is
45%
salt. She wants to obtain
110
ounces of a mixture that is
25%
salt. How many ounces of each solution should she use?
----
Equation:
salt + salt = salt
0.20x + 0.45(110-x) = 0.25*110
------------------
20x + 45*110 - 45x = 25*110
----
-25x = -10*110
----
x = (2/5)110
--
x = 2*22
x = 44 ounces (amt. of 20% solution needed)
110-44 = 66 ounces (amt. of 45% solution needed)
------------
Cheers,
Stan H.